The average of two numbers is 35. If the difference between them is 14, find the larger number?
Answer by Cromlix(4381)

Average of two numbers = 35
So, their total is 2*35 = 70
x + y = 70
x - y = 14
Add
2x = 84
x = 42
Substitute x = 42
in x + y = 70
42 + y = 70
y = 28
Larger number = 42.
Hope this helps.
:-)